OpenStudy (anonymous):

-8-(-4)divided by (-4)

OpenStudy (sasogeek):

-8+4/-4=-4/-4=1

OpenStudy (anonymous):

This is order of operations

OpenStudy (anonymous):

Can anybody help me with this?

OpenStudy (dumbcow):

2 negatives make a positive -8 -(-4) = -8 +4 = -4

OpenStudy (anonymous):

I need the answer.

OpenStudy (anonymous):

I don't know the answer.

OpenStudy (dumbcow):

haha...well sasogeek gave you that i showed you how to simplify the top, all you have to do is divide 2 numbers

OpenStudy (anonymous):

is the answer 4?

OpenStudy (dumbcow):

why do you think its 4? show your work

OpenStudy (anonymous):

idk

OpenStudy (dumbcow):

is this the problem? \[\frac{-8 -(-4)}{-4}\]

OpenStudy (anonymous):

I need the answer!

OpenStudy (dumbcow):

wow this is a forum to get help, not demand answers its simple arithmetic

OpenStudy (dumbcow):

i already told you that 2 negatives make a positive -8 -(-4) = -8 +4 = -4 \[\rightarrow \frac{-4}{-4}\] what does that equal?

OpenStudy (anonymous):



OpenStudy (dumbcow):

No, If i have 4 marbles and i have put them in 4 groups, how many marbles are in each group?

OpenStudy (anonymous):

1.

OpenStudy (dumbcow):

YES, the answer is 1 whenever you divide a number by itself you always get 1 if you noticed the very first post, sasogeek did it already

OpenStudy (anonymous):

ok